GlueX Level-3 Trigger Meeting, Apr 13, 2017

From GlueXWiki
Jump to: navigation, search

Meeting Info.

Meeting Time And Location

15:00 EST (JLab time)

CC A110


You can connect using BlueJeans Video conferencing (ID: 198 332 417). (Click "Expand" to the right for details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 198 332 417
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or
    • Enter meeting ID above

Background Information

(Click expand for some links to background information -->)


  1. Announcements
  2. Previous Meeting aka mini-review (Jul 22, 2016)
  3. Previously presented results
  4. Current Status
  5. May 18 review
  6. AOT


Attendees: Simon T., Elton S., Alex S., Dave A., Sergey F., David L.

  • Summary of final report from last July's L3 "mini-review" presented
  • Current status of DAQ parameters (as understood by David) was presented
    • DAQ event rate is currently limited by (according to Alex S.):
      1. CAEN TDCs
      2. TPOL
      3. CDC
    • 95kHz readout rate tested and should be used as one limiting parameter
    • L1 trigger goal according to Alex is 70kHz-80kHz
      • Currently unknown when accidentals become enough of an issue that increasing beam current does not gain us anything
      • This includes L1 rate reduction due to including ST in trigger. This sparked some discussion.
      • Current data rate estimates are up at or just over 2GB/s without ST. Closer to 1.5GB/s with ST
  • Two high intensity data sets were taken in Fall 2016. None were taken in Spring 2017./
  • Conversion from beam current to photon beam intensity on target not fully understood in David'd analysis
    • Alex described how he used PS rates and gets consistent results between Fall 2016 and Spring 2017 data
      • Transmission improved by roughly 20% since the Fall
  • We will meet again next week and weekly until the review
    • Action Items:
      1. Clarify beam intensity for Fall 2016 high intensity data sets and correlate to Spring 2017 conditions using PS (David)
      2. Estimate data volume and compare to what was previously provided to CC and use to estimate additional tape and hardware cost (David)
      3. Reinstall reprogrammed TD board in Hall (Alex)
      4. Continue testing of dual stream writing at high rate (Dave A., Sergey F.)